The 3505-kg anvil A of the drop forge is mounted on a nest of heavy coil springs having a combined stiffness of 2.2(106) N/m. The 590-kg hammer B falls 430 mm from rest and strikes the anvil, which suffers a maximum downward deflection of 20 mm from its equilibrium position. Determine the height h of rebound of the hammer and the coefficient of restitution e which applies.B 430 mm
Answer #1

Given M = 3505 kg stifiness ka 2-2x 10 N/m. hammer = 590kg, Falls from 430mm. Deflection = comm velocity of hammer = J2gh - vFrom Energy balance equation for anuit AE = AT + Over Aug 0 -4 mm, + } *x+ mgx-mgx - mut I kx (x= Deflection) = -1 * 3505x4,coefficient of estitution e = v-(-u) V e = 0.0663 - (-0.5) 2.904 = 0.195I forgot to upload 3rd image..please check the answer now..and take off the negative rating